# The Uniform Mordell–Lang Conjecture

### v3

@@ 3.1. Abelian variety
There are finitely many possibilities for the Hilbert polynomials of irreducible subvarieties of $A$ of dimension $r$ and degree $d$; see [Gro62, Thm.2.1(b) and Lem.2.4].[1][1][1]In the notation of [Gro62, Thm.2.1(b) and Lem.2.4], $X=A$, $S=\mathrm{Spec}\overline{\mathbb{Q}}$, $\mathcal{O}_{X}(1)=L$, $K=\overline{\mathbb{Q}}$, $Y$ the irreducible subvariety in question, $E$ the set of the classes of all the structural sheaves of such $Y$’s, and $\mathfrak{F}$ the structural sheaf of a such $Y$. The key point to prove this finiteness result is to compare the Hilbert scheme and the Chow variety. Let $\Xi$ be this finite set of polynomials.
Let $\mathbf{H}_{r,d}(A):=\bigcup_{P\in\Xi}\mathbf{H}_{P}(A)$, where $\mathbf{H}_{P}(A)$ is the Hilbert scheme which parametrizes subschemes of $A$ with Hilbert polynomial $P$. As $\Xi$ is a finite set, $\mathbf{H}_{r,d}(A)$ is of finite type over $\overline{\mathbb{Q}}$ and is projective.
There exists a universal family $\mathscr{X}_{r,d}(A)\rightarrow\mathbf{H}_{r,d}(A)$ endowed with a natural closed $\mathbf{H}_{r,d}(A)$-immersion
@@
where $\pi_{A}$ is the projection to the second factor. Over each point $s\in\mathbf{H}_{r,d}(A)(\overline{\mathbb{Q}})$, the fiber $\mathscr{X}_{r,d}(A)_{s}$ is precisely the subscheme of $A$ parametrized by $s$, and the horizontal immersion is the natural closed immersion $\mathscr{X}_{r,d}(A)_{s}$ in $A$.
By definition of Hilbert schemes, the morphism $\pi_{A}|_{\mathscr{X}_{r,d}(A)}$ is flat and proper. Define
(3.2) $\mathbf{H}_{r,d}(A)^{\circ}:=\{s\in\mathbf{H}_{r,d}(A)(\overline{\mathbb{Q}}):\mathscr{X}_{r,d}(A)_{s}\text{ is an integral subscheme of }A\}$
endowed with the reduced induced subscheme structure; hence it is a subvariety of $\mathbf{H}_{r,d}(A)$ which is quasi-projective over $\overline{\mathbb{Q}}$. Then $\mathbf{H}_{r,d}(A)^{\circ}$ parametrizes all irreducible subvarieties $X$ of $A$ with $\dim X=r$ and $\deg_{L}X=d$ (by our choice of $\Xi$ above). By [Gro67, Thm.12.2.4.(viii)], $\mathbf{H}_{r,d}(A)^{\circ}$ is Zariski open in $\mathbf{H}_{r,d}(A)$. For each irreducible component $V$ of $\mathbf{H}_{r,d}(A)$, the intersection $V\cap\mathbf{H}_{r,d}(A)^{\circ}$ is either Zariski open dense in $V$ or is empty.
@@
Let $V$ be a (not necessarily irreducible) subvariety of $\mathbf{H}_{r,d}(A)^{\circ}$. Then there exist $m_{0}$ points $P_{1},\ldots,P_{m_{0}}\in A(\overline{\mathbb{Q}})$ such that the Zariski closed subset of $V$ defined by
$\{[X]\in V(\overline{\mathbb{Q}}):P_{1}\in X(\overline{\mathbb{Q}}),~{}\ldots,P_{m_{0}}\in X(\overline{\mathbb{Q}})\}$
has dimension 00, i.e. this set is a non-empty finite set.
@@
Fix $[X_{0}]\in V(\overline{\mathbb{Q}})$. We claim for each $k\in\{0,\ldots,\dim V\}$ the following statement.
Claim There are finitely many points $P_{k,1},\ldots,P_{k,n_{k}}$ in $X_{0}(\overline{\mathbb{Q}})$ such that $V_{k}:=\{[X]\in V(\overline{\mathbb{Q}}):P_{k,1}\in X(\overline{\mathbb{Q}}),~{}\ldots,P_{k,n_{k}}\in X(\overline{\mathbb{Q}})\}$ has dimension $\leq\dim V-k$.
Notice that this claim immediately yields the lemma by taking $k=\dim V$; the set $V_{\dim V}$ thus obtained is non-empty since it contains $[X_{0}]$.
@@
Assume the claim holds true for $10,\ldots,k-$ 1. We have thus obtained point $P_{k-1,1},\ldots,$ $P_{k-1,n_{k-1}}\in X_{0}(\overline{\mathbb{Q}})$ such that $V_{k-1}:=\{[X]\in V(\overline{\mathbb{Q}}):P_{k-1,i}\in X(\overline{\mathbb{Q}})\text{ for all }i=1,\ldots,n_{k-1}\}$ has dimension $\leq\dim V-k+1$.
For each irreducible component $W$ of $V_{k-1}$ with $\dim W>0$, there exists some $[X_{W}]\in W(\overline{\mathbb{Q}})$ such that $X_{W}\not=X_{0}$ as (irreducible) subvarieties of $A$. Take $P_{W}\in(X_{0}\setminus X_{W})(\overline{\mathbb{Q}})$. Then $\{[X]\in W(\overline{\mathbb{Q}}):P_{W}\in X(\overline{\mathbb{Q}})\}$ has dimension $\leq\dim W-1\leq\dim V-k+1-1=\dim V-k$.
Thus it suffices to take $\{P_{k,1},\ldots,P_{k,n_{k}}\}:=\{P_{k-1,1},\ldots,$ $P_{k-1,n_{k-1}}\}\bigcup(\bigcup_{W}\{P_{W}\})$ with $W$ running over all positive dimensional irreducible components of $V_{k-1}$. ∎
@@
[…]
@@
Let $V$ be a (not necessarily irreducible) subvariety of $\mathbf{H}_{r,d}(A)^{\circ}$. Then there exists $m_{0}\geq 1$ with the following property. For each $m\geq m_{0}$, there exists $\underline{P}\in A^{m}(\overline{\mathbb{Q}})$ such that $(\iota_{A}^{[m]}|_{\mathscr{X}_{r,d}(A)^{[m]}\times_{\mathbf{H}_{r,d}(A)}V})^{-1}(\underline{P})$ has dimension 00, i.e. is a non-empty finite set.
@@
Let $P_{1},\ldots,P_{m_{0}}\in A(\overline{\mathbb{Q}})$ be from Lemma 3.1. For each $m\geq m_{0}$, set $P_{k}=P_{m_{0}}$ for each $k\geq m_{0}$.
@@
It is clear that $\pi_{A}^{[m]}|_{\{\underline{P}\}\times V}$ is an isomorphism. Therefore $\pi_{A}^{[m]}$ induces an isomorphism
($(\{\underline{P}\}\times V)\cap\mathscr{X}_{r,d}(A)^{[m]}\cong\pi_{A}^{[m]}\left((\{\underline{P}\}\times V)\cap\mathscr{X}_{r,d}(A)^{[m]}\right).$
Notice that the left hand side of this isomorphism is precisely $(\iota_{A}^{[m]}|_{\mathscr{X}_{r,d}(A)^{[m]}\times_{\mathbf{H}_{r,d}(A)}V})^{-1}(\underline{P})$. So it suffices to prove that the right hand side of this isomorphism is a non-empty finite set. A direct computation shows that the right hand side is
@@
which is non-empty finite by Lemma 3.1. Hence we are done. ∎
3.2. Family version
Let $\pi^{\mathrm{univ}}\colon\mathfrak{A}_{g}\rightarrow\mathbb{A}_{g}$ be the universal abelian variety over the fine moduli space of principally polarized abelian varieties of dimension $g$ with level-$4$-structure. For each $b\in\mathbb{A}_{g}(\overline{\mathbb{Q}})$, the abelian variety parametrized by $b$ is $(\mathfrak{A}_{g})_{b}=(\pi^{\mathrm{univ}})^{-1}(b)$.
Fix a symmetric relatively ample line bundle $\mathfrak{L}_{g}$ on $\mathfrak{A}_{g}/\mathbb{A}_{g}$ satisfying the following property: for each principally polarized abelian variety $(A,L)$ parametrized by $b\in\mathbb{A}_{g}(\overline{\mathbb{Q}})$, we have $c_{1}(\mathfrak{L}_{g}|_{(\mathfrak{A}_{g})_{b}})=2c_{1}(L)$; see [MFK94, Prop.6.10] for the existence of $\mathfrak{L}_{g}$.
Then $\mathfrak{L}_{g}^{\otimes 4}$ is relatively very ample on $\mathfrak{A}_{g}/\mathbb{A}_{g}$.
There are finitely many possibilities for the Hilbert polynomials of irreducible subvarieties of $(\mathfrak{A}_{g})_{b}$ (for all $b\in\mathbb{A}_{g}(\overline{\mathbb{Q}})$) of dimension $r$ and degree $d$ with respect to $\mathfrak{L}_{g}^{\otimes 4}|_{(\mathfrak{A}_{g})_{b}}$; see [Gro62, Thm.2.1(b) and Lem.2.4].[2][2][2]In the notation of [Gro62, Thm.2.1(b) and Lem.2.4], $X=\mathfrak{A}_{g}$, $S=\mathbb{A}_{g}$, $\mathcal{O}_{X}(1)=\mathfrak{L}_{g}^{\otimes 4}$, $K=\overline{\mathbb{Q}}$, $Y$ the irreducible subvariety in question, $E$ the set of the classes of all the structural sheaves of such $Y$’s, and $\mathfrak{F}$ the structural sheaf of a such $Y$. Let $\Xi$ be this finite set of polynomials.
Consider the Hilbert scheme
(3.4) $\mathbf{H}:=\bigcup_{P\in\Xi}\mathbf{H}_{P}(\mathfrak{A}_{g}/\mathbb{A}_{g})\xrightarrow{\iota_{\mathbf{H}}}\mathbb{A}_{g}$
with $\mathbf{H}_{P}(\mathfrak{A}_{g}/\mathbb{A}_{g})$ the $\mathbb{A}_{g}$-scheme representing the functor {schemes over $\{\text{schemes over }\mathbb{A}_{g}\}\rightarrow\{\text{sets}\}$, $T\mapsto\{\text{subschemes }W\text{ of }\mathfrak{A}_{g}\times_{\mathbb{A}_{g}}T\text{ which are proper flat over }T\text{ and have Hilbert polynomial }P\}$. It is known that each $\mathbf{H}_{P}(\mathfrak{A}_{g}/\mathbb{A}_{g})$ is a projective $\mathbb{A}_{g}$-scheme.
Then $\mathbf{H}$ is an $\mathbb{A}_{g}$-scheme with $\iota_{\mathbf{H}}$ the structural morphism. The morphism $\iota_{\mathbf{H}}$ is of finite type since $\Xi$ is a finite set.
There is a universal family endowed with a closed $\mathbf{H}$-immersion
(3.5) ${\mathscr{X}:=\mathscr{X}_{r,d}(\mathfrak{A}_{g}/\mathbb{A}_{g})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathfrak{A}_{g}\times_{\mathbb{A}_{g}}\mathbf{H}=:\mathcal{A}_{\mathbf{H}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi}$${\mathbf{H}}$
@@ 3.2. Family version
To ease notation, for each $\mathbf{H}$-scheme $\mathfrak{S}\rightarrow\mathbf{H}$ and for each morphism $S\rightarrow\mathbf{H}$, we will denote by $\mathfrak{S}_{S}:=\mathfrak{S}\times_{\mathbf{H}}S$. In particular, this applies to $\mathscr{X}$ and $\mathcal{A}_{\mathbf{H}}$; if $S$ is a variety, then $\mathcal{A}_{S}\rightarrow S$ is an abelian scheme, and $\mathscr{X}_{S}$ is a subvariety of $\mathcal{A}_{S}$ which dominates $S$.
As for (3.2), we define
(3.6) $\mathbf{H}^{\circ}=\mathbf{H}^{\circ}_{r,d}(\mathfrak{A}_{g}/\mathbb{A}_{g}):=\{s\in\mathbf{H}(\overline{\mathbb{Q}}):\mathscr{X}_{s}\text{ is an integral subscheme of }\mathcal{A}_{s}\}$
@@
endowed with the reduced induced subscheme structure; it is thus a quasi-projective variety defined over $\overline{\mathbb{Q}}$. We will call $\mathbf{H}^{\circ}$ the restricted Hilbert scheme.
@@
By our choice of $\Xi$ above (3.4), $\mathbf{H}^{\circ}(\overline{\mathbb{Q}})$ parametrizes all pairs $(X,(A,L))$ consisting of a principally polarized abelian variety defined over $\overline{\mathbb{Q}}$ and an irreducible subvariety $X$ of $A$ defined over $\overline{\mathbb{Q}}$ with $\dim X=r\geq 1$ and $\deg_{\mathfrak{L}_{g}^{\otimes 4}|_{A}}X=d$.[3][3][3]Here $(A,L)$ gives rise to a point $b\in\mathbb{A}_{g}(\overline{\mathbb{Q}})$, and we identify $A$ with $(\mathfrak{A}_{g})_{b}$. By [Gro67, Thm.12.2.4.(viii)] applied to $\pi|_{\mathscr{X}}$, $\mathbf{H}^{\circ}$ is Zariski open in $\mathbf{H}$. For each irreducible component $S$ of $\mathbf{H}$, the intersection $S\cap\mathbf{H}^{\circ}$ is either Zariski open dense in $S$ or is empty. Moreover for the structural morphism $\iota_{\mathbf{H}}\colon\mathbf{H}\rightarrow\mathbb{A}_{g}$, we have that $\iota_{\mathbf{H}}|_{\mathbf{H}^{\circ}}^{-1}(b)$ is precisely the $\mathbf{H}_{r,d}(A_{b})^{\circ}$ defined in (3.2).
Write $\iota_{\mathbf{H}^{\circ}}:=\iota_{\mathbf{H}}|_{\mathbf{H}^{\circ}}$.
Convention. From now on, we will forget the scheme $\mathbf{H}$ and only work with the variety $\mathbf{H}^{\circ}$ defined over $\overline{\mathbb{Q}}$; notice that $\mathbf{H}^{\circ}$ is in general not irreducible.
@@
For each $m\geq 1$, set $\mathscr{X}_{\mathbf{H}^{\circ}}^{[m]}$ (resp. $\mathcal{A}_{\mathbf{H}^{\circ}}^{[m]}$) to be the $m$-fibered power of $\mathscr{X}_{\mathbf{H}^{\circ}}$ over $\mathbf{H}^{\circ}$ (resp. of $\mathcal{A}_{\mathbf{H}^{\circ}}$ over $\mathbf{H}^{\circ}$). Then we have a commutative diagram
@@
(3.7) ${\mathscr{X}^{[m]}_{\mathbf{H}^{\circ}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{A}^{[m]}_{\mathbf{H}^{\circ}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi^{[m]}}$$\scriptstyle{\iota^{[m]}}$${\mathfrak{A}_{g}^{[m]}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi^{\mathrm{univ},[m]}}$${\mathbf{H}^{\circ}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\iota_{\mathbf{H}^{\circ}}}$${\mathbb{A}_{g}}$
Over each $b\in\mathbb{A}_{g}(\overline{\mathbb{Q}})$, this diagram restricts to (3.3) (with $A=(\mathfrak{A}_{g})_{b}$) completed by $(\mathfrak{A}_{g})_{b}^{m}\rightarrow\{b\}$ and $\mathbf{H}_{r,d}((\mathfrak{A}_{g})_{b})^{\circ}\rightarrow\{b\}$.
@@
Let $S$ be an irreducible subvariety of $\mathbf{H}^{\circ}$. Then there exists $m_{0}\geq 1$ with the following property. For each $m\geq m_{0}$, $\iota^{[m]}|_{\mathscr{X}^{[m]}_{S}}$ is generically finite.

### v4

@@ 3.1. Parameterizing subvarieties of an abelian variety
The following result can be deduced from [Gro62, Thm. 2.1(b) and Lem. 2.4], which itself is proved by comparing the Hilbert scheme and the Chow variety: There is a finite set $\Xi$ of polynomials such that the Hilbert polynomial of any irreducible subvariety of $A$ of dimension $r$ and degree $d$ is an element of $\Xi$. Indeed, in the notation of [Gro62, Thm. 2.1(b) and Lem. 2.4], we can take $X=A$, $S=\mathrm{Spec}(\overline{\mathbb{Q}})$, $\mathcal{O}_{X}(1)=L$, $K=\overline{\mathbb{Q}}$, $Y\subseteq A$ the irreducible subvarieties of degree $d$, $\mathfrak{F}$ the structure sheaf of such a subvariety $Y$, and $E$ the set of the classes of all these structure sheaves.
Let $\mathbf{H}_{r,d}(A):=\bigcup_{P\in\Xi}\mathbf{H}_{P}(A)$, where $\mathbf{H}_{P}(A)$ is the Hilbert scheme that parameterizes subschemes of $A$ with Hilbert polynomial $P$ and is constructed in [Gro62, Thm. 3.2]. As $\Xi$ is a finite set, $\mathbf{H}_{r,d}(A)$ is a projective variety over $\overline{\mathbb{Q}}$. By construction, there exists a universal family $\mathscr{X}_{r,d}(A)\rightarrow\mathbf{H}_{r,d}(A)$ endowed with a natural closed $\mathbf{H}_{r,d}(A)$-immersion
@@
where $\pi_{A}$ is the projection to the second factor. Over each point $s\in\mathbf{H}_{r,d}(A)(\overline{\mathbb{Q}})$, the fiber $\mathscr{X}_{r,d}(A)_{s}$ is precisely the subscheme of $A$ parametrized by $s$, and the horizontal immersion restricts to its natural closed immersion in $A$.
By the definition of Hilbert schemes, the morphism $\pi_{A}|_{\mathscr{X}_{r,d}(A)}$ is flat and proper. By [Gro67, Thm. 12.2.4.(viii)],
(3.2) $\{s\in\mathbf{H}_{r,d}(A):\mathscr{X}_{r,d}(A)_{s}\text{ is geometrically integral}\}$
is a Zariski open subset of $\mathbf{H}_{r,d}(A)$, which we endow with the induced subscheme structure. We write $\mathbf{H}_{r,d}(A)^{\circ}$ for the maximal reduced subscheme of this scheme. Note that $\mathbf{H}_{r,d}(A)^{\circ}$ is a quasi-projective variety defined over $\overline{\mathbb{Q}}$. Its points $\mathbf{H}_{r,d}(A)^{\circ}(\overline{\mathbb{Q}})$ parametrize all irreducible $\overline{\mathbb{Q}}$-subvarieties $X$ of $A$ with $\dim X=r$ and $\deg_{L}X=d$ (by our choice of $\Xi$ above). For each irreducible component $V$ of $\mathbf{H}_{r,d}(A)$, the intersection $V\cap\mathbf{H}_{r,d}(A)^{\circ}$ is either a dense Zariski open in $V$ or empty.
@@
Let $V$ be a (not necessarily irreducible) subvariety of $\mathbf{H}_{r,d}(A)^{\circ}$. Then there exist $m_{0}=m_{0}(V)$ points $P_{1},\ldots,P_{m_{0}}\in A(\overline{\mathbb{Q}})$ such that the Zariski closed subset of $V$ defined by
$\{[X]\in V(\overline{\mathbb{Q}}):P_{1}\in X(\overline{\mathbb{Q}}),\ldots,P_{m_{0}}\in X(\overline{\mathbb{Q}})\}$
has dimension 0, i.e. is a non-empty finite set.
@@
Fix $X_{0}]\in V(\overline{\mathbb{Q}})$. For each $k\in\{0,\ldots,\dim V\}$, we claim the following statement: There are finitely many points $P_{k,1},\ldots,P_{k,n_{k}}$ in $X_{0}(\overline{\mathbb{Q}})$ such that
$V_{k}:=\{[X]\in V(\overline{\mathbb{Q}}):P_{k,1}\in X(\overline{\mathbb{Q}}),\ldots,P_{k,n_{k}}\in X(\overline{\mathbb{Q}})\}$
has dimension $\leq\dim V-k$.
Taking $k=\dim V$, this claim immediately yields the lemma; the set $V_{\dim V}$ thus obtained is non-empty since it contains $X_{0}]$.
@@
Assume the claim holds true for $0,\ldots,k-1$. We have thus obtained points $P_{k-1,1},\ldots,$ $P_{k-1,n_{k-1}}\in X_{0}(\overline{\mathbb{Q}})$ such that
$V_{k-1}:=\{[X]\in V(\overline{\mathbb{Q}}):P_{k-1,i}\in X(\overline{\mathbb{Q}})\text{ for all }i=1,\ldots,n_{k-1}\}$
has dimension $\leq\dim V-k+1$.
For each irreducible component $W$ of $V_{k-1}$ with $\dim W>0$, there exists some $X_{W}]\in W(\overline{\mathbb{Q}})$ such that $X_{W}\not=X_{0}$ as (irreducible) subvarieties of $A$. Take $P_{W}\in(X_{0}\setminus X_{W})(\overline{\mathbb{Q}})$. Then $\{[X]\in W(\overline{\mathbb{Q}}):P_{W}\in X(\overline{\mathbb{Q}})\}$ has dimension $\leq\dim W-1\leq\dim V-k+1-1=\dim V-k$.
Thus it suffices to take $\{P_{k,1},\ldots,P_{k,n_{k}}\}:=\{P_{k-1,1},\ldots,$ $P_{k-1,n_{k-1}}\}\cup(\bigcup_{W}\{P_{W}\})$ with $W$ running over all positive dimensional irreducible components of $V_{k-1}$. ∎
@@
${\mathscr{X}_{r,d}(A)^{[m]}\times_{\mathbf{H}_{r,d}(A)}\mathbf{H}_{r,d}(A)^{\circ}\ \ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${A^{m}\times\mathbf{H}_{r,d}(A)^{\circ}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi_{A}^{[m]}}$$\scriptstyle{\iota_{A}^{[m]}}$${A^{m}}$${\mathbf{H}_{r,d}(A)^{\circ}}$
@@
Let $V$ be a (not necessarily irreducible) subvariety of $\mathbf{H}_{r,d}(A)^{\circ}$. Then there exists $m_{0}=m_{0}(V)\geq 1$ with the following property. For each $m\geq m_{0}$, there exists $\underline{P}\in A^{m}(\overline{\mathbb{Q}})$ such that
$(\iota_{A}^{[m]}|_{\mathscr{X}_{r,d}(A)^{[m]}\times_{\mathbf{H}_{r,d}(A)}V})^{-1}(\underline{P})$
has dimension 0, i.e. is a non-empty finite set.
@@
Let $P_{1},\ldots,P_{m_{0}}\in A(\overline{\mathbb{Q}})$ be from Lemma 3.1. For each $k\geq m_{0}$, we set $P_{k}=P_{m_{0}}$.
@@
It is clear that $\pi_{A}^{[m]}|_{\{\underline{P}\}\times V}$ is an isomorphism. Therefore, $\pi_{A}^{[m]}$ induces an isomorphism
$(\{\underline{P}\}\times V)\cap\mathscr{X}_{r,d}(A)^{[m]}\cong\pi_{A}^{[m]}\left((\{\underline{P}\}\times V)\cap\mathscr{X}_{r,d}(A)^{[m]}\right).$
The left hand side of this isomorphism is precisely $(\iota_{A}^{[m]}|_{\mathscr{X}_{r,d}(A)^{[m]}\times_{\mathbf{H}_{r,d}(A)}V})^{-1}(\underline{P})$. So it suffices to prove that the right hand side of this isomorphism is a non-empty finite set. A direct computation shows that the right hand side is
@@
which is non-empty and finite by Lemma 3.1. ∎
3.2. Parameterizing subvarieties in families
We next extend the setting of the previous subsection to the case of families of abelian varieties. Let $\pi^{\mathrm{univ}}\colon\mathfrak{A}_{g}\rightarrow\mathbb{A}_{g}$ be the universal abelian variety over the fine moduli space of principally polarized abelian varieties of dimension $g$ with level-$4$-structure. For each $b\in\mathbb{A}_{g}(\overline{\mathbb{Q}})$, the abelian variety parametrized by $b$ is $(\mathfrak{A}_{g})_{b}=(\pi^{\mathrm{univ}})^{-1}(b)$.
We fix a symmetric relatively ample line bundle $\mathfrak{L}_{g}$ on $\mathfrak{A}_{g}/\mathbb{A}_{g}$ satisfying the following property: for each principally polarized abelian variety $(A,L)$ parametrized by $b\in\mathbb{A}_{g}(\overline{\mathbb{Q}})$, we have $c_{1}(\mathfrak{L}_{g}|_{(\mathfrak{A}_{g})_{b}})=2c_{1}(L)$; see [MFK94, Prop. 6.10] for the existence of $\mathfrak{L}_{g}$. The line bundle $\mathfrak{L}_{g}^{\otimes 4}$ is relatively very ample on $\mathfrak{A}_{g}/\mathbb{A}_{g}$.
Again by [Gro62, Thm. 2.1(b) and Lem. 2.4], there are finitely many possibilities for the Hilbert polynomials of irreducible subvarieties of $(\mathfrak{A}_{g})_{b}$ (for all $b\in\mathbb{A}_{g}(\overline{\mathbb{Q}})$) of dimension $r$ and degree $d$ with respect to $\mathfrak{L}_{g}^{\otimes 4}|_{(\mathfrak{A}_{g})_{b}}$.[1][1][1]In the notation of [Gro62, Thm. 2.1(b) and Lem. 2.4], $X=\mathfrak{A}_{g}$, $S=\mathbb{A}_{g}$, $\mathcal{O}_{X}(1)=\mathfrak{L}_{g}^{\otimes 4}$, $K=\overline{\mathbb{Q}}$, $Y$ the irreducible subvarieties of degree $d$, $\mathfrak{F}$ the structure sheaf of such a subvariety $Y$, and $E$ the set of the classes of all these structure sheaves. Let $\Xi$ be this finite set of polynomials.
Consider the $\mathbb{A}_{g}$-scheme
(3.4) $\mathbf{H}:=\bigsqcup_{P\in\Xi}\mathbf{H}_{P}(\mathfrak{A}_{g}/\mathbb{A}_{g})\xrightarrow{\iota_{\mathbf{H}}}\mathbb{A}_{g}$
with $\mathbf{H}_{P}(\mathfrak{A}_{g}/\mathbb{A}_{g})$ the $\mathbb{A}_{g}$-scheme representing the functor from $\mathbb{A}_{g}$-schemes to sets associating with an $\mathbb{A}_{g}$-scheme $T$ the set of proper subschemes $W\subseteq\mathfrak{A}_{g}\times_{\mathbb{A}_{g}}T$ that are flat over $T$ and have Hilbert polynomial $P$. Each $\mathbf{H}_{P}(\mathfrak{A}_{g}/\mathbb{A}_{g})$ is projective over $\mathbb{A}_{g}$. Since $\Xi$ is a finite set, the structure morphism $\iota_{\mathbf{H}}$ is of finite type.
Similarly to the previous subsection, there is a universal family endowed with a closed $\mathbf{H}$-immersion
(3.5) ${\mathscr{X}:=\mathscr{X}_{r,d}(\mathfrak{A}_{g}/\mathbb{A}_{g})\ \ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathfrak{A}_{g}\times_{\mathbb{A}_{g}}\mathbf{H}=:\mathcal{A}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi}$${\mathbf{H}}$
@@ 3.2. Parameterizing subvarieties in families
To ease notation, we set $\mathfrak{S}_{S}:=\mathfrak{S}\times_{\mathbf{H}}S$ for every $\mathbf{H}$-scheme $\mathfrak{S}\rightarrow\mathbf{H}$ and every morphism $S\rightarrow\mathbf{H}$. In particular, we do so for $\mathscr{X}$ and $\mathcal{A}$ as well as for their fibered self-products over $\mathbf{H}$. If $S$ is a variety, then $\mathcal{A}_{S}\rightarrow S$ is an abelian scheme, and $\mathscr{X}_{S}$ is a subvariety of $\mathcal{A}_{S}$ which dominates $S$.
As for (3.2), the set
(3.6) $\{s\in\mathbf{H}:\mathscr{X}_{r,d}(\mathfrak{A}_{g}/\mathbb{A}_{g})_{s}\text{ is geometrically integral}\}$
@@
is Zariski open in $\mathbf{H}$ and hence has an induced structure as an open subscheme. We write $\mathbf{H}_{r,d}(\mathfrak{A}_{g}/\mathbb{A}_{g})^{\circ}$ for the maximal reduced subscheme of this subscheme, which is a quasi-projective variety defined over $\overline{\mathbb{Q}}$. In the sequel, $\mathbf{H}_{r,d}(\mathfrak{A}_{g}/\mathbb{A}_{g})^{\circ}$ is called the restricted Hilbert scheme. In addition, we denote it simply by $\mathbf{H}^{\circ}$ if $r$, $d$ and $g$ are clear from context.
@@
By our choice of $\Xi$ and the construction (3.4), $\mathbf{H}^{\circ}(\overline{\mathbb{Q}})$ parametrizes all pairs $((A,L),X)$ consisting of a principally polarized abelian variety defined over $\overline{\mathbb{Q}}$ and an irreducible subvariety $X$ of $A$ defined over $\overline{\mathbb{Q}}$ with $\dim X=r\geq 1$ and $\deg_{\mathfrak{L}_{g}^{\otimes 4}|_{A}}X=d$.[2][2][2]Here $(A,L)$ gives rise to a point $b\in\mathbb{A}_{g}(\overline{\mathbb{Q}})$, and we identify $A$ with $(\mathfrak{A}_{g})_{b}$. For each irreducible component $S$ of $\mathbf{H}$, the intersection $S\cap\mathbf{H}^{\circ}$ is either a dense Zariski open set in $S$ or empty. For convenience, we set $\iota_{\mathbf{H}^{\circ}}:=\iota_{\mathbf{H}}|_{\mathbf{H}^{\circ}}$ for the restriction of the structural morphism. The fiber $\iota_{\mathbf{H}^{\circ}}^{-1}(b)$ is precisely the subvariety $\mathbf{H}_{r,d}(A_{b})^{\circ}$ defined in the previous subsection.
@@
For each $m\geq 1$, let $\mathscr{X}_{\mathbf{H}^{\circ}}^{[m]}$ (resp. $\mathcal{A}_{\mathbf{H}^{\circ}}^{[m]}$) denote the $m$-fold fibered power of $\mathscr{X}_{\mathbf{H}^{\circ}}$ over $\mathbf{H}^{\circ}$ (resp. of $\mathcal{A}_{\mathbf{H}^{\circ}}$ over $\mathbf{H}^{\circ}$). There is a commutative diagram
@@
(3.7) ${\mathscr{X}^{[m]}_{\mathbf{H}^{\circ}}\ \ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{A}^{[m]}_{\mathbf{H}^{\circ}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi^{[m]}}$$\scriptstyle{\iota^{[m]}}$${\mathfrak{A}_{g}^{[m]}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi^{\mathrm{univ},[m]}}$${\mathbf{H}^{\circ}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\iota_{\mathbf{H}^{\circ}}}$${\mathbb{A}_{g}.}$
Taking fibers in this diagram over a point $b\in\mathbb{A}_{g}(\overline{\mathbb{Q}})$, we obtain again the diagram (3.3) for the abelian variety $A=(\mathfrak{A}_{g})_{b}$.
@@
Let $S$ be an irreducible subvariety of $\mathbf{H}^{\circ}$. Then there exists $m_{0}=m_{0}(S)\geq 1$ such that the restriction of $\iota^{[m]}$ to $\mathscr{X}_{S}^{[m]}$ is generically finite onto its image for every $m\geq m_{0}$.